OpenStudy (kitkat16):

Help Please!! CD s perpendicular to AB and passes through point C(5, 12). If the coordinates of A and B are (-10, -3) and (7, 14), respectively, the x- intercept of CD is(12,0),(15,0),(17,0),(19,0) The point (-5,24)(-2,19)(7,-10)(8,11) lies on CD

OpenStudy (dopeness_):

What's the question? What are you trying to find?

Directrix (directrix):

Please post the question as it appears in your text. Thanks.

OpenStudy (anonymous):



OpenStudy (anonymous):

The rest doesnt really make sense

OpenStudy (kitkat16):

Its a two part question. A and B are (-10, -3) and (7, 14), the x- intercept of CD is ? these are my choices: (12,0),(15,0),(17,0),(19,0) The second part of the question is : What point lies on CD? these are my choices: (-5,24),(-2,19),(7,-10),(8,11)

OpenStudy (kitkat16):

When I draw it out on graph paper the only one I see working would be points (19,0) and point (-2, 19). if you graph a line in passes through point (5,12), is this correct?

OpenStudy (kitkat16):

Is there an equation I can do to fiqure this out? I'm pretty sure there is but I can't remember.

Directrix (directrix):

>>>the x- intercept of CD I don't know what this means. If you could post the question exactly as it appears in your text, it would be helpful.

OpenStudy (kitkat16):

I did at the top when I posted the question

OpenStudy (kitkat16):

ok give me a min

OpenStudy (kitkat16):

@dayakar can you help me please?

OpenStudy (dayakar):

find the slope of a line AB = m1 = (y2-y1)/ (x2-x1) can u do that?

OpenStudy (kitkat16):

ok give me a min

OpenStudy (kitkat16):

the slope is 1

OpenStudy (dayakar):

what happens?

OpenStudy (dayakar):

very good

OpenStudy (kitkat16):

m=1

OpenStudy (dayakar):

slope of a line CD which is perpendicular to AB = m2 = 1/m1 = -1 is it clear

OpenStudy (dayakar):

m2 = -1/m1 = -1

OpenStudy (kitkat16):

so I have to find the slope of 1 for CD

OpenStudy (dayakar):

do u understand

OpenStudy (dayakar):

now we going to find equation of a line CD, whose slope (m2) =-1 and which is passing through C(5,12) =(x1, y1) is y-y1 = m2 (x-x1) can u write equation

OpenStudy (kitkat16):

ok give me a min

OpenStudy (kitkat16):

m=(19-(0)/(-2-(17)

OpenStudy (kitkat16):

(17,0) and (-2,19)

OpenStudy (dayakar):

there is no need of find m we know m2 = -1

OpenStudy (dayakar):

y- y1 = m2 (x-x1) [ m2=-1, (x1,y1) = (5,12) ]

OpenStudy (dayakar):

y- 12 = -1 (x-5)

OpenStudy (dayakar):

can u simplify it

OpenStudy (kitkat16):

x=17

OpenStudy (dayakar):

y-12 = -x +5 x+y = 12+5 x+y = 17-------(1) if u understand this we can easily find the answers

OpenStudy (dayakar):

do u understand or not

OpenStudy (kitkat16):

no I'm sorry

OpenStudy (dayakar):

what is ur doubt

OpenStudy (kitkat16):

I understand your equation above but don't understand how to apply it

OpenStudy (dayakar):

1) to find x- intercept plug y=0 in equation (1) now what is the x value?

OpenStudy (kitkat16):

-2

OpenStudy (dayakar):

x+y =17 ----(1) x+0 =17 [ since y=0 ] x= 17 required point = (x ,y ) = (17 ,0)

OpenStudy (dayakar):

is it clear

OpenStudy (kitkat16):

yes I got 17,0. I thought i was trying to figure out where the point lies

OpenStudy (dayakar):



OpenStudy (dayakar):

can we move to second one

OpenStudy (kitkat16):

yes please

OpenStudy (dayakar):

sorry it is not C , i want to draw rough diagram where the point (17,0) lie

OpenStudy (dayakar):

2) it is very simple, do u have any idea

OpenStudy (kitkat16):

no sorry

OpenStudy (dayakar):

x+y =17 -----(1) check all the options add x and y which equals to 17 that is the answer

OpenStudy (kitkat16):

-2,19

OpenStudy (kitkat16):

-2+19 =17

OpenStudy (dayakar):

excellent , that's it

OpenStudy (kitkat16):

omgosh thank you

OpenStudy (dayakar):

:)

OpenStudy (kitkat16):

i made myself crazy with this

OpenStudy (dayakar):

my graph is wrong remove it , it is not important here

OpenStudy (dayakar):

once again go through it , u can get the concept good luck

OpenStudy (kitkat16):

I can't remove graph I think only you can. Thanks so much for helping me.

OpenStudy (kitkat16):

did someone else give you the medal? I went to give you one and it was already clicked. Crazy. lol

OpenStudy (kitkat16):

I guess it was me thanks again!!!